40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 2 16:29:40 2020, gautam, Update, Electronics, c1psl VME crate removed, Acro-crate installed c1psl.pdf
    Reply  Tue Mar 3 17:20:14 2020, gautam, Update, Electronics, More cabling removed 
       Reply  Wed Mar 4 16:18:31 2020, gautam, Update, Electronics, More cabling removed 1X2EuroBefore.JPGIOO.png
          Reply  Wed Mar 4 21:02:49 2020, Koji, Update, Electronics, More cabling removed 
    Reply  Tue Mar 3 17:59:33 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
       Reply  Thu Mar 5 15:03:48 2020, Yehonathan, Update, Electronics, PSL Shutter and PMC TRANSPD working 
          Reply  Thu Mar 5 19:45:23 2020, Jon, Summary, PSL, C1PSL in-situ test results c1psl_feedthrough_wiring_-_By_Connector_(3).pdf
             Reply  Wed Mar 11 18:12:53 2020, gautam, Summary, PSL, WFS Demod board modifications D980233-B_Mar2020Mods.pdf
Message ID: 15255     Entry time: Thu Mar 5 15:03:48 2020     In reply to: 15243     Reply to this: 15256
Author: Yehonathan 
Type: Update 
Category: Electronics 
Subject: PSL Shutter and PMC TRANSPD working 

[Jon, Yehonathan]

Summary

With the Acromag chassis now permanently installed, we tested the C1PSL channels going over the channel list one by one, excluding the IMC channels which Gautam is taking responsibility for (the servo board itself is also in question).

The strategy is to check the response of input channels to specific output channels for expected behaviour whenever is possible.

We marked on the channel list spreadsheet the status of the channels that were tested.

In more detail

FSS

Channels under test What was done
C1:PSL-FSS_SW1 Switched C1:PSL-FSS_SW1 and observed the IMC unlock
C1:PSL-FSS_SW2, C1:PSL-FSS_MIXERM Connected a signal to Test2 on FSS box and observed a proportional change on C1:PSL-FSS_MIXERM
C1:PSL-FSS_INOFFSET Disconnected feedback by switching C1:PSL-FSS_SW1. Tweaked C1:PSL-FSS_INOFFSET and observed a proportional response in C1:PSL-FSS_MIXERM
C1:PSL-FSS_MGAIN, C1:PSL-FSS_PCDRIVE Disconnected feedback, turned on some offset using C1:PSL-FSS_INOFFSET. Tweaked C1:PSL-FSS_MGAIN and observed a response in C1:PSL-FSS_PCDRIVE
C1:PSL-FSS_SLOWDC, C1:PSL-FSS_SLOWM Disconnected feedback. Tweaked C1:PSL-FSS_SLOWDC and obsereved a proportional response in C1:PSL-FSS_SLOWM
C1:PSL-FSS_FASTGAIN, C1:PSL-FSS_FAST Disconnected feedback, turned on some offset using C1:PSL-FSS_INOFFSET. Tweaked C1:PSL-FSS_FASTGAIN and obsereved a response in  C1:PSL-FSS_FAST

 

Frequency Ref

Channels under test What was done
C1:PSL-PMC_PHCON Observed the PMC unlocks when a big change in C1: PSL-PMC_PHCON is made
C1:PSL-PMC_RFADJ, C1:PSL-PMC_MODET Tweaked C1:PSL-PMC_RFADJ and obsereved a proportional response in C1:PSL-PMC_MODET
C1:PSL-PMC_PHFLIP Observed the PMC unlock when C1:PSL-PMC_PHFLIP is switched

 

PMC Servo Card

Channels under test What was done
C1:PSL-PMC_SW1, C1:PSL-PMC_PMCERR, C1:PSL-PMC_INOFFSET, C1:PSL-PMC_PZT Unlocked the PMC by switching C1:PSL-PMC_SW1. Tweaked C1:PSL-PMC_INOFFSET and observed a proportional change in C1:PSL-PMC_PMCERR and C1:PSL-PMC_PZT
C1:PSL-PMC_BLANK Observed the PMC unlock with when C1:PSL-PMC_BLANK is switched
C1:PSL-PMC_GAIN Unlocked the PMC by switching C1:PSL-PMC_SW1. Turned on some offset using  C1:PSL-PMC_INOFFSET. Tweaked C1:PSL-PMC_GAIN and observed response in C1:PSL-PMC_PZT
C1:PSL-PMC_SW2 Unlocked the PMC by switching C1:PSL-PMC_SW1. Connected a signal to TP2 on the PMC card and observed a proportional change in C1:PSL-PMC_PZT.
C1:PSL-PMC_RAMP

Unlocked the PMC by switching C1:PSL-PMC_SW1. Tweaked C1:PSL-PMC_RAMP and observed a change in C1:PSL-PMC_PZT.

C1:PSL-PMC_RFPDDC Observed a high value 0.5V when PMC is unlocked and a low value 0.03V when it is locked

 

WFSs

Channels under test What was done
C1:IOO-WFS*_SEG*_ATTEN

We misaligned MC1 to get a measurable signal in WFS channels. NDScoped the corresponding C1:IOO-WFS*_SEG*_I&Q channels and observed a change in those channels in response to switching the attenuation on and off.

C1:IOO-WFS*_LO_LOCK_MON Disconnected the LO cable from the WFS boards and observed C1:IOO-WFS*_LO_LOCK_MON go to zero.
C1:IOO-WFS*_SEG*_I&Q Connected a short SMA cable to the 29.5MHz frequency distribution board. Attenuated the signal by 20db and connected it to the different SEG channels one at a time and observed a response in C1:IOO-WFS*_SEG*_I&Q channels.
C1:IOO-WFS*_SEG*_DC We shined a laser pointer to the different quadrants and observed saturation in the corresponding C1:IOO-WFS*_SEG*_DC with no cross talks.

MC Servo

Channels under test What was done
C1:IOO-MC_SW1, C1:IOO-MC_OPTIONA, C1:IOO-MC_POL, C1:IOO-MC_OPTIONB,C1:IOO-MC_FASTSW These switches unlocked the IMC when flipped.
C1:IOO-MC_SW2, C1:IOO-MC_SUM_MON, C1:IOO-MC_SLOW_MON, C1:IOO-MC_FAST_MON A sine wave signal was injected in IN2 on the servo board. C1:IOO-MC_SW2 was switched on and the value of C1:IOO-MC_SUM_MON, C1:IOO-MC_SLOW_MON and C1:IOO-MC_FAST_MON changed accordingly.
C1:IOO-MC_SW3 Connected a scope to OUT2 on the servo board. Switched C1:IOO-MC_SW3 on and observed a signal on the scope.
C1:IOO-MC_EXCA_EN Unlocked the IMC by switching C1:IOO-MC_SW1 off. Connected a signal to EXC A and a scope to TP2A on the servo board and observed the signal on the scope when C1:IOO-MC_EXCA_EN was switched on.
C1:IOO-MC_EXCB_EN Unlocked the IMC by switching C1:IOO-MC_SW1 off. Connected a signal to EXC B and a scope to TP2B on the servo board and observed the signal on the scope when C1:IOO-MC_EXCB_EN was switched on.
C1:IOO-MC_REFL_OFFSET Unlocked the IMC by switching off. Tweaked C1:IOO-MC_REFL_OFFSET and observed a proportional change in C1:IOO-MC_SUM_MON.
C1:IOO-MC_LATCH_EN Tweaked the VCO gain slider and observed the latch switch off and on.
C1:IOO-MC_LIMIT Unlocked the IMC by switching C1:IOO-MC_SW1 off. Connected a sine wave signal to EXC B and enabled C1:IOO-MC_EXCB_EN. Ramped up the VCO gain. Raised the sine wave amplitude until C1:IOO-MC_LIMIT turned on.
C1:IOO-MC_LIMITER We ramped the VCO such that C1:IOO-MC_LIMIT was switched on. We switched C1:IOO-MC_LIMITER on and observed C1:PSL-FSS_MIXERM high value go down.

NPRO Diagnostics

Channels under test What was done
C1:PSL-NPRO_*

The signals were compared to previous values for consistency. Then they were unplugged from the Acromag chassis to confirm their values went to 0 and returned to the same values after being reconnected.

ELOG V3.1.3-